A cell undergoing meiosis has four pairs of homologous chromosomes. How many different combinations of maternal and paternal chromosomes are possible within the daughter cells that form? (Do not consider crossing over in your answer.)
a.16
b. 8
c. 4
d. 2
Answer #1

Number of chromosome combinations in crossing over = 2^n, where n is haploid chromosome number

Here, n = 4

2n = 8

Number of chromosome combinations = 2^4 = 16

Answer is 16.

Please rate high.